The following data represent a random sample of 15 marks out of 20) on a Statistics quiz. Assume that the marks are normally distributed 4 7 7 5 3 10 6 8 7 9 8 9 5 4 7 a. Determine the standard deviation of the marks. b. Estimate the population mean with 95% confidence.


1
Expert's answer
2022-04-13T16:25:33-0400

First let's find mean:

"\mu=\frac{1}{15}(4+7+7+5+3+10+6+8+\\
+7+9+8+9+5+4+7)=\\
=\frac{1}{15}(3+2\cdot4+2\cdot5+6+\\
4\cdot7+2\cdot8+2\cdot9+10)=6.6"


a) Variance:

"\sigma^2=\frac{1}{n-1}\sum_{i=1}^{n}(x_i-\mu)^2=\\
=\frac{1}{14}(3.6^2+2\cdot2.6^2+2\cdot1.6^2+0.6^2+\\
+4\cdot0.4^2+2\cdot1.4^2+2\cdot2.4^2+3.4^2)\approx4.257"

Standard deviation:

"\sigma=\sqrt{\sigma^2}\approx2.063"


b) t-value for 15 - 1 = 14 degrees of freedom and 95% confidence is 2.145

Confidence interval:

"\mu\pm t\frac{\sigma}{\sqrt{n}}=6.6\pm 2.145\frac{2.063}{\sqrt{15}}=6.60\pm 1.14"